OpenStudy (sanaelshazly):

A parallel plate capacitor is connected to a battery for charging. after some time, while the battery is still connected to the capacitor, the distance between the capacitor plates is doubled. which of the following statements are true ? a) the charge on the plates does not change. b) the capacitance doubles. c) the potential difference of the battery is halved. d) the electric field between the plates is halved. e) the potential difference across the plates does not change. f) none of the above.

OpenStudy (sanaelshazly):

what does that mean !

OpenStudy (experimentx):

when d increase, capacitance decrease ...

OpenStudy (sanaelshazly):

so !! which answer is the right one !

OpenStudy (experimentx):

you guess!!

OpenStudy (sanaelshazly):

f !!

OpenStudy (sanaelshazly):

or D ! right !

OpenStudy (sanaelshazly):

i think it's d !

OpenStudy (experimentx):

it don't think D

OpenStudy (sanaelshazly):

what !

OpenStudy (sanaelshazly):

so which answer !

OpenStudy (experimentx):

guess

OpenStudy (sanaelshazly):

so it's f !

OpenStudy (experimentx):

there might be some correct.

OpenStudy (anonymous):

its D.

OpenStudy (experimentx):

might be ... but i think e might be correct too.

OpenStudy (sanaelshazly):

hey ! i want one answer for god sake :D u both are driving me crazy !

OpenStudy (experimentx):

i haven't checked the possibility of d) Oo.. V = Ed => E = V/d => E = V/2d guess this is right too.

OpenStudy (anonymous):

V is directly proportional to Q and the increase in d will cause the P.D to decrease as v=kq/r so D is the correct one

OpenStudy (experimentx):

still, it seems the potential across the place is same.

OpenStudy (experimentx):

LOL ... i got confused, you can do with D

OpenStudy (sanaelshazly):

:D ! finally :D

OpenStudy (experimentx):

well sorry for the mess.

OpenStudy (sanaelshazly):

no it's ok :D ! ThnQQ !

OpenStudy (anonymous):

@experimentX , the electric field is sigma/ (epsilon nought) between 2 plates of capacitor. How does it change ??

OpenStudy (anonymous):

According to me the answer is none of above are correct.

OpenStudy (anonymous):

Because capacitance becomes half Potential becomes twice

OpenStudy (anonymous):

Charge is constant

OpenStudy (anonymous):

electric field remains constant

OpenStudy (experimentx):

my initial guess was ... the potential remains constant.  as the capacitor is charged, there is no current ... hence the pd must be equal to emf.